OpenStudy (anonymous):

inverse laplace transform of 3s+5/(s^2-6s+25)

OpenStudy (anonymous):

Complete the square in the denominator first: \[\begin{align*}s^2-6s+25&=s^2-6s+9+16\\ &=(s-3)^2+16 \end{align*}\] \[\begin{align*}F(s)&=\frac{3s+5}{(s-3)^2+16}\\ &=\frac{3s-9+14}{(s-3)^2+16}\\ &=\frac{3(s-3)}{(s-3)^2+16}+\frac{14}{(s-3)^2+16}\\ \mathcal{L}^{-1}\left\{F(s)\right\}&=3\mathcal{L}^{-1}\left\{\frac{s-3}{(s-3)^2+16}\right\}+\frac{14}{4}\mathcal{L}^{-1}\left\{\frac{4}{(s-3)^2+16}\right\}\\ f(t)&=3e^{3t}\cos4t+\frac{14}{4}e^{3t}\sin4t \end{align*}\]

OpenStudy (anonymous):

A "missing" step in red, added to clarify where the \(e^{3t}\) comes from. \[\begin{align*}F(s)&=\frac{3s+5}{(s-3)^2+16}\\ &=\frac{3s-9+14}{(s-3)^2+16}\\ &=\frac{3(s-3)}{(s-3)^2+16}+\frac{14}{(s-3)^2+16}\\ \mathcal{L}^{-1}\left\{F(s)\right\}&=3\mathcal{L}^{-1}\left\{\frac{s-3}{(s-3)^2+16}\right\}+\frac{14}{4}\mathcal{L}^{-1}\left\{\frac{4}{(s-3)^2+16}\right\}\\ &=\color{red}{3e^{3t}\mathcal{L}^{-1}\left\{\frac{s}{s^2+16}\right\}+\frac{14}{4}e^{3t}\mathcal{L}^{-1}\left\{\frac{4}{s^2+16}\right\}}\\ f(t)&=3e^{3t}\cos4t+\frac{14}{4}e^{3t}\sin4t \end{align*}\]